Given the function f(x)=4x^2+8x-2. Calculate the following values:
f(x+1)=
f(-x)=

Answer by stanbon(75887)

Given the function f(x)=4x^2+8x-2.
Calculate the following values:
f(x+1)= 4(x+1)^2 + 8(x+1) - 2 = 4(x^2+2x+1) + 8x+8 -2 = 4x^2 +16x + 10
------------------------------
f(-x)= 4(-x)^2 + 8(-x) - 2 = 4x^2 -8x -2
-------------
Cheers,
Stan H.
------------